# Structure and Independence in Hyperbolic Uniform Disk Graphs

### v1

@@ 5.3 Candidate nooses and noose hierarchies
Proof.
Let $S$ be a independent set of $G$ with $|S|\leq k$. As the vertices of $S$ have pairwise distance at least $r$, the Delaunay complex $\mathcal{D}(S)$ of $S$ is $1+\mathcal{O}\big{(}\frac{\log k}{r}\big{)}$-outerplanar by Theorem 2. The outerplanarity of a graph gives a linear upper bound on its treewidth [Bodlaender98, Theorem 83] and thus also its branchwidth [Robertson_Seymour_1991, (5.1)]. Thus, $\mathcal{D}(S)$ has branchwidth $w\in\mathcal{O}\big{(}1+\frac{\log k}{r}\big{)}$, i.e., each noose of any minimum width sphere cut decomposition of $\mathcal{D}(S)$ visits at most $w$ sites. As candidate nooses are required to be nooses in minimum width branch decompositions, each candidate noose visits at most $w$ vertices. There are at most $n^{w}$ different sequences in which a candidate noose can visit at most $w$ vertices of $G$.
@@
Observe that the above estimate is constructive, i.e., we can efficiently enumerate all possible options and filter out sequences of points that do not give a generalized polygon (e.g., as it would be self-intersecting). Note that not all options actually yield valid candidate nooses. However, we clearly get a superset of $\mathcal{C}$ in the desired time. ∎
@@
Now consider any independent set $S$ of $G$ with Delaunay complex $\mathcal{D}(S)$. Each noose of a minimum width normalized sphere cut decomposition of $\mathcal{D}(S)$ is a generalized polygon and the nooses of such a sphere cut decomposition form a polygon hierarchy. Observe that the above definition of being valid is directly derived from the properties stated in Lemma 5.2 and thus this hierarchy is valid. Moreover, due to Lemma 5.2, it is also well-spaced. Note that the statement of Lemma 5.2 is actually stronger, as it guarantees the distance requirements globally and not only locally for every child–parent relation. Thus, the sphere cut decomposition of $\mathcal{D}(S)$ yields a valid and well-spaced polygon hierarchy, leading directly to the following corollary. {corollary} Let $G$ be a hyperbolic uniform disk graph with radius $r$ and let $S$ be an independent set of $G$. Then the nooses of a normalized sphere cut decomposition of $\mathcal{D}(S)$ form a valid and well-spaced polygon hierarchy of $G$ whose generalized polygons visit all vertices of $S$.
@@
In the first case, $v$ is also visited by $O_{L}$ or one if its descendants. Thus $u,v\in S(O_{L})$ and they have pairwise distance at least $2r$ by induction. In the second case, $v$ lies on $O_{R}$ but not on $O_{L}$ and thus in the exterior of $O_{L}$. As the hierarchy is well-spaced, $v$ has distance at least $r$ from $O_{L}$. Moreover, by induction, $u$ lies in the interior of $O_{L}$ and has distance at least $r$ from $O_{L}$. Thus, $u$ and $v$ have distance at least $2r$. For the third and final case, $v$ lies in the interior of $O_{R}$ and has distance at least $r$ from $O_{R}$ by induction. As the interiors of $O_{L}$ and $O_{R}$ are disjoint, the line segment between $u$ and $v$ has to cover the distance of at least $r$ from $u$ to $O_{L}$ and the distance of at least $r$ from $O_{R}$ to $v$. Thus, also in this final case, $u$ and $v$ have distance at least $2r$. ∎
@@ 5.4 Solving independent set
At the same time Independent Set in $G\in\mathrm{HUDG}(r)$ can also be decided in $2^{\mathcal{O}(\sqrt{n})}$ time or $n^{\mathcal{O}(\sqrt{k})}$ time. To see this, recall that the uniformly sized hyperbolic disks representing the vertices of $G$ can also be viewed as Euclidean disks in the Poincaré disk. This means that $G$ is also an intersection graph of $n$ disks in the Euclidean plane. As shown by de Berg et al. [BergBKMZ20, Corollary 2.4] and by Marx and Pilipczuk [MarxP22], Independent Set in a disk graph can be decided in $2^{\mathcal{O}(\sqrt{n})}$ time or $n^{\mathcal{O}(\sqrt{k})}$ time. ∎
@@ 6 Independent set approximation
Our approximation algorithm in Theorem 4 works similarly to Lipton and Tarjan’s version for planar graphs [LiptonT80]: we repeatedly apply a separator until we get small patches and then solve each patch individually using the algorithm of Theorem 3. Unlike in planar graphs, we do not have a strong and readily available lower bound on the size of a maximum independent set to inform us how much of the graph we can afford to ignore, so we will start by proving one based on degeneracy.
@@
Figure 11: Figures for the proof of Lemma 6.
Draw a horizontal line through $c$ and two lines at angle $\pi/3$ to form six wedges. The lower three wedges cannot contain any disk centers, because they would be further from the origin than $c$. Now, shrink each disk $D^{\prime}$ intersecting $D$ until it has the same radius as $D$, while keeping the point of $D^{\prime}$ closest to $c$ fixed; see Figure 11(a). The resulting disk will be a subset of $D^{\prime}$ and have its center in the same wedge. This makes the situation as in Figure 11(b) and means that we can use the same arguments as for unit disks [Peeters91]: for each wedge, the disks with center in that wedge that intersect $D$ must form a clique with $D$, giving three cliques in total.
Now, we will stab $D$ and the disks it intersects. For this, assume without loss of generality that $D$ has unit radius, then consider the problem of using unit disks to cover the upper half of a disk of radius $2$ centered at $c$. This can be done with four disks, as shown in Figure 11(c). The centers of these covering disks are our stabbing points: any disk $\overline{D}$ intersecting $D$ must have its center in the half-disk and thus $\overline{D}$ will be stabbed by the center of the covering disk that contains the center of $\overline{D}$. ∎
By repeatedly picking the vertex given by Lemma 6, the clique bound gives a simple 3-approximation algorithm for Independent Set. This generalizes the algorithm for unit disks by Marathe et al. [MaratheBHRR95]. Additionally, we get that any hyperbolic uniform disk graph is $(3\omega-3)$-degenerate with $\omega$ being the size of the largest clique. It is also $(4\ell-1)$-degenerate for ply $\ell$, so the maximum independent set has size $\Omega(n/\ell)$.
As first step in the approximation algorithm, we will use the separators of De Berg et al. [BergBKMZ20] to separate the graph into many small patches [^11] For certain $r$ it will be better to use the separator of Theorem 1, but this only improves constants in exponents.. These separators are $36/37$-balanced, have size $\mathcal{O}(\sqrt{n})$, and can be found in $\mathcal{O}(n^{4})$ time. The following proof matches that of Frederickson [Frederickson87] that is the main lemma for so-called $r$-divisions [^12] We are forced to use $t$ because $r$ already refers to the disk radius..
{lemma}
@@
Proof.
We will iteratively apply a balanced clique-based separator until the size of the graph drops below $t$ vertices. Let $B(n,t)$ denote the number of cliques we need to remove from a graph of size $n$ and target size $t$. If $n\leq t$, nothing needs to be done, so $B(n,t)=0$. Otherwise, use the separator of De Berg et al. [BergBKMZ20] to separate $G$ into two induced subgraphs of roughly equal size and recurse on these. Let $c$ be such that the number of cliques in the separator is always at most $c\sqrt{n}$. We will use induction to prove that $B(n,t)\leq 43cn/\sqrt{t}-7c\sqrt{n}$ for $n>t/37$. As base case we have $t/37<n\leq t$; here we established $B(n,t)=0$ and we have $\sqrt{37n/t}>1$, which implies $43cn/\sqrt{t}>7c\sqrt{n}$ so the bound holds. Now, for $n>t$ we get
@@
$\leq c\sqrt{n}+43cn/\sqrt{t}-7c\sqrt{n}\cdot\left(\sqrt{1/37}+\sqrt{36/37}\right)$
@@
$\leq 43cn/\sqrt{t}-7c\sqrt{n}.$
Thus, $B(n,t)\in\mathcal{O}(n/\sqrt{t})$. The recursion has depth $\mathcal{O}(\log(n/t))$ and at each level we find separators in disjoint subgraphs of $G$. Thus, this procedure takes at most $\mathcal{O}(n^{4})$ time per level and $\mathcal{O}(n^{4}\log(n/t))$ time in total. ∎
This gives us all we need to prove Theorem 4. See 4
Proof.
Let $t=\ell^{2}/\varepsilon^{2}$ and apply Lemma 6, taking $\mathcal{O}(n^{4}\log n)$ time and removing $\mathcal{O}(\varepsilon n/\ell)$ cliques. Each of these cliques could only have contributed one point to the independent set. Since we know the maximum independent set has size $\Omega(n/\ell)$, we can only have thrown away a $\mathcal{O}(\varepsilon)$ fraction of it, so finding the exact maximum independent set in the remaining graph gives a $(1-\mathcal{O}(\varepsilon))$-approximation. We do this by applying Theorem 3 separately to each of the $\mathcal{O}(n/t)$ subgraphs of size $t$. This takes $t^{\mathcal{O}(1+\frac{\log t}{r})}=(\frac{\ell}{\varepsilon})^{\mathcal{O}(1+\frac{1}{r}\log\frac{\ell}{\varepsilon})}$ time for each subgraph and thus $n\cdot(\frac{\ell}{\varepsilon})^{\mathcal{O}(1+\frac{1}{r}\log\frac{\ell}{\varepsilon})}$ time in total. This makes the total running time $\mathcal{O}\left(n^{4}\log n\right)+n\cdot(\frac{\ell}{\varepsilon})^{\mathcal{O}(1+\frac{1}{r}\log\frac{\ell}{\varepsilon})}$ to find a $(1-\mathcal{O}(\varepsilon))$-approximation, so by appropriately adjusting $\varepsilon$ we can also get a $(1-\varepsilon)$-approximation with the same asymptotic running time. ∎
Using the results of Har-Peled [HarPeled23], Theorem 4 also directly implies a similarly efficient approximation algorithm for Minimum Vertex Cover. A set $S$ of vertices in $G$ forms a vertex cover if every edge of $G$ is incident to some vertex of $S$. In the Minimum Vertex Cover problem we are given a graph $G$ and $\varepsilon>0$, and we wish to output that a vertex cover $S$ of $G$ that has size $(1+\varepsilon)k^{*}$, where $k^{*}$ is the size of the minimum vertex cover.
{corollary}
@@ 7 Algorithmic consequences of our separator
We can directly use our clique-based separator to compute independent sets and prove Corollary 1 via a simple divide-and conquer algorithm as in Corollary 2.4 of [BergBKMZ20]. In our case, we use the trivial bound of $n$ on the size of the cliques in Theorem 1, which means that the weight of each clique in the separator is at most $\log(n+1)$. Consequently, the separator’s weight is at most $\mathcal{O}((1+1/r)\log^{2}n)$. The proof of Corollary 1 now follows that of Corollary 2.4 of [BergBKMZ20], we merely replace the separator weight $\mathcal{O}(n^{1-1/d})$ with $\mathcal{O}((1+1/r)\log^{2}n)$. We give the proof for completeness.
See 1
Proof.
The algorithm works as follows: First, we find a balanced separator according to Theorem 1 in $\mathcal{O}(n\log n)$ time. We can then iterate over all possible intersections $X$ between a fixed optimum independent set and the separator. For every clique we can pick one of at most $n$ vertices, or nothing, so there are $(n+1)^{\mathcal{O}(1+1/r)\log n}=2^{\mathcal{O}(1+1/r)\log^{2}n}$ options. For each of these possible intersections, we remove all disks intersecting $X$, and recurse on both sides of the separator.
@@
This concludes the proof. ∎
We note that the recursion added an extra logarithmic factor, which can be avoided via $\mathcal{P}$-flattened treewidth. We do not optimize this running time as the resulting algorithm would still be slower than that of Theorem 3. However, we do need to bound the $\mathcal{P}$-flattened treewidth for Corollary 1.
Consider a graph $G\in\mathrm{HUDG}(r)$ where $r\in\mathcal{O}(1)$, let $C$ be any clique of $G$, and let $s$ be the center of some disk of $C$. Then all other disk centers from $C$ are contained in a disk of radius $2r$ around $s$. Since $r\in\mathcal{O}(1)$, a disk of radius $r$ has area $\Theta(r^{2})$, and the above argument shows that the union of the disks in $C$ occupy area $\Theta(r^{2})$. We will now decompose $\mathbb{H}^{2}$ into regions of diameter at most $2r$ such that each region has an inscribed disk of radius at least $\Omega(r)$. The following can be derived from Lemma 2.1 (ii) of [Kisfaludi-Bak20] and from Theorem 6 of [Kisfaludi-BakW24].
{lemma}
[see [Kisfaludi-Bak20, Kisfaludi-BakW24]] For any $\delta\in\mathcal{O}(1)$ there exists a subdivision of the hyperbolic plane into connected regions with the following properties:
@@
We are now ready to prove Corollary 1. The proof merely needs to establish a clique partition $\mathcal{P}$ with a corresponding weighted treewidth bound, and the property that the graph $G_{\mathcal{P}}$ has maximum constant degree, that is, each clique of $\mathcal{P}$ has at most constantly many neighboring cliques. Then the machinery of [BergBKMZ20] yields the desired algorithms automatically.
See 1
Proof.
Use the subdivision of Lemma 7 with $\delta=2r$. We can define a clique-partition $\mathcal{P}$ of $G$ by putting disks in the same partition class whenever their centers fall in the same region of the subdivision. Clearly each created class induces a clique in $G$ as the disk centers have pairwise distance less or equal to the diameter of the region, which is at most $2r$.
Fix such a clique partition $\mathcal{P}$, and consider now a separator from Theorem 1. If $X$ is one of the cliques in the separator, then let $p_{x}$ be the center of some disk of $X$. If $C\in\mathcal{P}$ and $X$ contain disks that intersect each other, then the graph distance of any pair of disks in $C$ and $X$ is at most $3$. Consequently, all disk centers in $C\cup X$ are covered by a disk $D$ of radius $3\cdot 2r=6r$ around $p_{x}$. Since each region of $\mathcal{P}$ has area $\Omega(r^{2})$, there can be at most $\mathcal{O}(1)$ regions of the subdivision contained inside $D$. In particular, $X$ can intersect at most $\mathcal{O}(1)$ cliques from $\mathcal{P}$.
Consider all cliques from $\mathcal{P}$ that are intersected by the separator cliques of Theorem 1: they clearly form a separator, and the above argument shows that there are $\mathcal{O}((1+1/r)\log n)$ such cliques in $\mathcal{P}$. Since each clique of $\mathcal{P}$ has weight at most $1+\log n$, we get a separator of weight $\mathcal{O}((1+1/r)\log^{2}n)$ over $\mathcal{P}$. Now [BergBKMZ20] (see also [Kisfaludi-Bak20]) implies that the $\mathcal{P}$-flattened treewidth of $G$ is $\mathcal{O}((1+1/r)\log^{2}n)$. Since $\mathcal{P}$ is a clique-partition with the additional property that each partition class is neighboring to $\mathcal{O}(1)$ other cliques, we can apply the entire machinery of [BergBKMZ20] and derive the required algorithms. ∎

### v2

@@ 7.3 Candidate nooses and noose hierarchies
Proof 7.8.
Let $S$ be a independent set of $G$ with $|S|\leq k$. As the vertices of $S$ have pairwise distance at least $r$, the Delaunay complex $\mathcal{D}(S)$ of $S$ is $1+\mathcal{O}\big{(}\frac{\log k}{r}\big{)}$-outerplanar by Theorem 1.4. The outerplanarity of a graph gives a linear upper bound on its treewidth [10, Theorem 83] and thus also its branchwidth [43, (5.1)]. Thus, $\mathcal{D}(S)$ has branchwidth $w\in\mathcal{O}\big{(}1+\frac{\log k}{r}\big{)}$, i.e., each noose of any minimum width sphere cut decomposition of $\mathcal{D}(S)$ visits at most $w$ sites. As candidate nooses are required to be nooses in minimum width branch decompositions, each candidate noose visits at most $w$ vertices. Each visited vertex is one of $n$ vertices of $G$ and thus there are at most $n^{w}$ different sequences in which a candidate noose can visit at most $w$ vertices of $G$.
@@
Observe that the above estimate is constructive, i.e., we can efficiently enumerate all possible options and filter out sequences of points that do not give a generalized polygon (e.g., as it would be self-intersecting). Note that not all options actually yield valid candidate nooses. However, we clearly get a superset of $\mathcal{C}$ in the desired time.
@@
Now consider any independent set $S$ of $G$ with Delaunay complex $\mathcal{D}(S)$. Each noose of a minimum width normalized sphere cut decomposition of $\mathcal{D}(S)$ is a generalized polygon and the nooses of such a sphere cut decomposition form a polygon hierarchy. Observe that the above definition of being valid is directly derived from the properties stated in Lemma 7.3 and thus this hierarchy is valid. Moreover, due to Lemma 7.5, it is also well-spaced. Note that the statement of Lemma 7.5 is actually stronger, as it guarantees the distance requirements globally and not only locally for every child–parent relation. Thus, the sphere cut decomposition of $\mathcal{D}(S)$ yields a valid and well-spaced polygon hierarchy, leading directly to the following corollary.
Corollary 7.9.
Let $G$ be a hyperbolic uniform disk graph with radius $r$ and let $S$ be an independent set of $G$. Then the nooses of a normalized sphere cut decomposition of $\mathcal{D}(S)$ form a valid and well-spaced polygon hierarchy of $G$ whose generalized polygons visit all vertices of $S$.
@@
In the first case, $v$ is also visited by $O_{L}$ or one if its descendants. Thus $u,v\in S(O_{L})$ and they have pairwise distance at least $2r$ by induction. In the second case, $v$ lies on $O_{R}$ but not on $O_{L}$ and thus in the exterior of $O_{L}$. As the hierarchy is well-spaced, $v$ has distance at least $r$ from $O_{L}$. Moreover, by induction, $u$ lies in the interior of $O_{L}$ and has distance at least $r$ from $O_{L}$. Thus, $u$ and $v$ have distance at least $2r$. For the third and final case, $v$ lies in the interior of $O_{R}$ and has distance at least $r$ from $O_{R}$ by induction. As the interiors of $O_{L}$ and $O_{R}$ are disjoint, the line segment between $u$ and $v$ has to cover the distance of at least $r$ from $u$ to $O_{L}$ and the distance of at least $r$ from $O_{R}$ to $v$. Thus, also in this final case, $u$ and $v$ have distance at least $2r$.
@@ 7.4 Solving independent set
At the same time Independent Set in $G\in\mathrm{HUDG}(r)$ can also be decided in $2^{\mathcal{O}(\sqrt{n})}$ time or $n^{\mathcal{O}(\sqrt{k})}$ time. To see this, recall that the uniformly sized hyperbolic disks representing the vertices of $G$ can also be viewed as Euclidean disks in the Poincaré disk. This means that $G$ is also an intersection graph of $n$ disks in the Euclidean plane. As shown by De Berg, Bodlaender, Kisfaludi-Bak, Marx, and Van der Zanden [15, Corollary 2.4] and by Marx and Pilipczuk [37], Independent Set in a disk graph can be decided in $2^{\mathcal{O}(\sqrt{n})}$ time or $n^{\mathcal{O}(\sqrt{k})}$ time.
@@ 8 Independent set approximation
Our approximation algorithm in Theorem 1.6 works similarly to Lipton and Tarjan’s version for planar graphs [33]: we repeatedly apply a separator until we get small patches and then solve each patch individually using the algorithm of Theorem 1.5. Unlike in planar graphs, we do not have a strong and readily available lower bound on the size of a maximum independent set to inform us how much of the graph we can afford to ignore, so we will start by proving one based on degeneracy.
@@
Figure 13: Figures for the proof of Lemma 8.1.
Draw a horizontal line through $c$ and two lines at angle $\pi/3$ to form six wedges. The lower three wedges cannot contain any disk centers, because they would be further from the origin than $c$. Now, shrink each disk $D^{\prime}$ intersecting $D$ until it has the same radius as $D$, while keeping the point of $D^{\prime}$ closest to $c$ fixed; see Figure 13(a). The resulting disk will be a subset of $D^{\prime}$ and have its center in the same wedge. This makes the situation as in Figure 13(b) and means that we can use the same arguments as for unit disks [41]: for each wedge, the disks with center in that wedge that intersect $D$ must form a clique with $D$, giving three cliques in total.
Now, we will stab $D$ and the disks it intersects. For this, assume without loss of generality that $D$ has unit radius, then consider the problem of using unit disks to cover the upper half of a disk of radius $2$ centered at $c$. This can be done with four disks, as shown in Figure 13(c). The centers of these covering disks are our stabbing points: any disk $\overline{D}$ intersecting $D$ must have its center in the half-disk and thus $\overline{D}$ will be stabbed by the center of the covering disk that contains the center of $\overline{D}$.
By repeatedly picking the vertex given by Lemma 8.1, the clique bound gives a simple 3-approximation algorithm for Independent Set. This generalizes the algorithm for unit disks by Marathe, Breu, Hunt III, Ravi, and Rosenkrantz [34]. Additionally, we get that any hyperbolic uniform disk graph is $(3\omega-3)$-degenerate with $\omega$ being the size of the largest clique. It is also $(4\ell-1)$-degenerate for ply $\ell$, so the maximum independent set has size $\Omega(n/\ell)$.
As first step in the approximation algorithm, we will use the separators of De Berg, Bodlaender, Kisfaludi-Bak, Marx, and Van der Zanden [15] to separate the graph into many small patches [^13] For certain $r$ it will be better to use the separator of Theorem 1.1, but this only improves constants in exponents.. These separators are $36/37$-balanced, have size $\mathcal{O}(\sqrt{n})$, and can be found in $\mathcal{O}(n^{4})$ time. The following proof matches that of Frederickson [18] that is the main lemma for so-called $r$-divisions [^14] We are forced to use $t$ because $r$ already refers to the disk radius..
Lemma 8.3.
@@
Proof 8.4.
We will iteratively apply a balanced clique-based separator until the size of the graph drops below $t$ vertices. In particular, we use the separator of De Berg, Bodlaender, Kisfaludi-Bak, Marx, and Van der Zanden [15], which for a disk graph of $n$ vertices consists of up to $c\sqrt{n}$ cliques, for some constant $c$. Let $B(n,t)$ denote the number of cliques we need to remove from a graph of size $n$ and target size $t$; we will use induction to prove that $B(n,t)<43cn/\sqrt{t}-7c\sqrt{n}$ for $n>t/37$. As base case we have $t/37<n\leq t$. Seeing as $n\leq t$, nothing needs to be done, so $B(n,t)=0$. Simultaneously, $n>t/37$ implies $43cn/\sqrt{t}>\frac{43}{\sqrt{47}}c\sqrt{n}>7c\sqrt{n}$, so indeed $B(n,t)=0<43cn/\sqrt{t}-7c\sqrt{n}$.
If $n>t$, we use the separator to get two induced subgraphs that are both guaranteed to have size at least $n/37$, then recurse on these. This gives the following:
@@
$<c\sqrt{n}+43cn/\sqrt{t}-7c\sqrt{\frac{n}{37}}-7c\sqrt{\frac{36n}{37}}$
@@
$<43cn/\sqrt{t}-7c\sqrt{n}.$
Thus, $B(n,t)\in\mathcal{O}(n/\sqrt{t})$. The recursion has depth $\mathcal{O}(\log(n/t))$ and at each level we find separators in disjoint subgraphs of $G$. Thus, this procedure takes at most $\mathcal{O}(n^{4})$ time per level and $\mathcal{O}(n^{4}\log(n/t))$ time in total.
This gives us all we need to prove Theorem 1.6. See 1.6
Proof 8.5.
Let $t=\ell^{2}/\varepsilon^{2}$ and apply Lemma 8.3, taking $\mathcal{O}(n^{4}\log n)$ time and removing $\mathcal{O}(\varepsilon n/\ell)$ cliques. Each of these cliques could only have contributed one point to the independent set. Since we know the maximum independent set has size $\Omega(n/\ell)$, we can only have thrown away a $\mathcal{O}(\varepsilon)$ fraction of it, so finding the exact maximum independent set in the remaining graph gives a $(1-\mathcal{O}(\varepsilon))$-approximation. We do this by applying Theorem 1.5 separately to each of the $\mathcal{O}(n/t)$ subgraphs of size $t$. This takes $t^{\mathcal{O}(1+\frac{\log t}{r})}=(\frac{\ell}{\varepsilon})^{\mathcal{O}(1+\frac{1}{r}\log\frac{\ell}{\varepsilon})}$ time for each subgraph and thus $n\cdot(\frac{\ell}{\varepsilon})^{\mathcal{O}(1+\frac{1}{r}\log\frac{\ell}{\varepsilon})}$ time in total. This makes the total running time $\mathcal{O}\left(n^{4}\log n\right)+n\cdot(\frac{\ell}{\varepsilon})^{\mathcal{O}(1+\frac{1}{r}\log\frac{\ell}{\varepsilon})}$ to find a $(1-\mathcal{O}(\varepsilon))$-approximation, so by appropriately adjusting $\varepsilon$ we can also get a $(1-\varepsilon)$-approximation with the same asymptotic running time.
Using the results of Har-Peled [23], Theorem 1.6 also directly implies a similarly efficient approximation algorithm for Minimum Vertex Cover. A set $S$ of vertices in $G$ forms a vertex cover if every edge of $G$ is incident to some vertex of $S$. In the Minimum Vertex Cover problem we are given a graph $G$ and $\varepsilon>0$, and we wish to output that a vertex cover $S$ of $G$ that has size $(1+\varepsilon)k^{*}$, where $k^{*}$ is the size of the minimum vertex cover.
Corollary 8.6.
@@ 9 Algorithmic consequences of our separator
We can directly use our clique-based separator to compute independent sets and prove Corollary 1.2 via a simple divide-and conquer algorithm as in Corollary 2.4 of [15]. In our case, we use the trivial bound of $n$ on the size of the cliques in Theorem 1.1, which means that the weight of each clique in the separator is at most $\log(n+1)$. Consequently, the separator’s weight is at most $\mathcal{O}((1+1/r)\log^{2}n)$. The proof of Corollary 1.2 now follows that of Corollary 2.4 of [15], we merely replace the separator weight $\mathcal{O}(n^{1-1/d})$ with $\mathcal{O}((1+1/r)\log^{2}n)$. We give the proof for completeness.
See 1.2
Proof 9.1.
The algorithm works as follows: First, we find a balanced separator according to Theorem 1.1 in $\mathcal{O}(n\log n)$ time. We can then iterate over all possible intersections $X$ between a fixed optimum independent set and the separator. For every clique we can pick one of at most $n$ vertices, or nothing, so there are $(n+1)^{\mathcal{O}(1+1/r)\log n}=2^{\mathcal{O}(1+1/r)\log^{2}n}$ options. For each of these possible intersections, we remove all disks intersecting $X$, and recurse on both sides of the separator.
@@
This concludes the proof.
We note that the recursion added an extra logarithmic factor, which can be avoided via $\mathcal{P}$-flattened treewidth. We do not optimize this running time as the resulting algorithm would still be slower than that of Theorem 1.5. However, we do need to bound the $\mathcal{P}$-flattened treewidth for Corollary 1.3.
Consider a graph $G\in\mathrm{HUDG}(r)$ where $r\in\mathcal{O}(1)$, let $C$ be any clique of $G$, and let $s$ be the center of some disk of $C$. Then all other disk centers from $C$ are contained in a disk of radius $2r$ around $s$. Since $r\in\mathcal{O}(1)$, a disk of radius $r$ has area $\Theta(r^{2})$, and the above argument shows that the union of the disks in $C$ occupy area $\Theta(r^{2})$. We will now decompose $\mathbb{H}^{2}$ into regions of diameter at most $2r$ such that each region has an inscribed disk of radius at least $\Omega(r)$. The following can be derived from Lemma 2.1 (ii) of [27] and from Theorem 6 of [29].
Lemma 9.2 (see [27, 29]).
For any $\delta\in\mathcal{O}(1)$ there exists a subdivision of the hyperbolic plane into connected regions with the following properties:
@@
We are now ready to prove Corollary 1.3. The proof merely needs to establish a clique partition $\mathcal{P}$ with a corresponding weighted treewidth bound, and the property that the graph $G_{\mathcal{P}}$ has maximum constant degree, that is, each clique of $\mathcal{P}$ has at most constantly many neighboring cliques. Then the machinery of [15] yields the desired algorithms automatically.
See 1.3
Proof 9.3.
Use the subdivision of Lemma 9.2 with $\delta=2r$. We can define a clique-partition $\mathcal{P}$ of $G$ by putting disks in the same partition class whenever their centers fall in the same region of the subdivision. Clearly each created class induces a clique in $G$ as the disk centers have pairwise distance less or equal to the diameter of the region, which is at most $2r$.
Fix such a clique partition $\mathcal{P}$, and consider now a separator from Theorem 1.1. If $X$ is one of the cliques in the separator, then let $p_{x}$ be the center of some disk of $X$. If $C\in\mathcal{P}$ and $X$ contain disks that intersect each other, then the graph distance of any pair of disks in $C$ and $X$ is at most $3$. Consequently, all disk centers in $C\cup X$ are covered by a disk $D$ of radius $3\cdot 2r=6r$ around $p_{x}$. Since each region of $\mathcal{P}$ has area $\Omega(r^{2})$, there can be at most $\mathcal{O}(1)$ regions of the subdivision contained inside $D$. In particular, $X$ can intersect at most $\mathcal{O}(1)$ cliques from $\mathcal{P}$.
Consider all cliques from $\mathcal{P}$ that are intersected by the separator cliques of Theorem 1.1: they clearly form a separator, and the above argument shows that there are $\mathcal{O}((1+1/r)\log n)$ such cliques in $\mathcal{P}$. Since each clique of $\mathcal{P}$ has weight at most $1+\log n$, we get a separator of weight $\mathcal{O}((1+1/r)\log^{2}n)$ over $\mathcal{P}$. Now [15] (see also [27]) implies that the $\mathcal{P}$-flattened treewidth of $G$ is $\mathcal{O}((1+1/r)\log^{2}n)$. Since $\mathcal{P}$ is a clique-partition with the additional property that each partition class is neighboring to $\mathcal{O}(1)$ other cliques, we can apply the entire machinery of [15] and derive the required algorithms.